5x - 2y = -5
y - 5x = 3

Answer by tutorcecilia(2144)

5x - 2y = -5 [Solve using substitution]
y - 5x = 3
.
y-5x=3 [Pick one equation and solve for one of the variables]
y=5x+3 [Substitute back into the other equation]
.
5x - 2(5x+3) = -5 [Simplify]
5x-10x-6=-5
-5x=-5+6 [Solve for x]
-5x=1
x=-1/5
.
y - 5x = 3 [Plug (x=-1/5)
y-5(-1/5)=3
y+1=3 [Solve for y]
y=-1+3
y=2
.
Check by plugging (x=-1/5) and (y=2) into the other equation.